# Diffusion Models for Computational Design at the Example of Floor Plans

### v1

@@ Abstract
AI Image generators based on diffusion models are widely discussed recently for their capability to create images from simple text prompts. But, for practical use in civil engineering they need to be able to create specific construction plans for given constraints. Within this paper we explore the capabilities of those diffusion-based AI generators for computational design at the example of floor plans and identify their current limitation. We explain how the diffusion-models work and propose new diffusion models with improved semantic encoding. In several experiments we show that we can improve validity of generated floor plans from 6 % to 90 % and query performance for different examples. We identify short comings and derive future research challenges of those models and discuss the need to combine diffusion models with building information modelling. With this we provide key insights into the current state and future directions for diffusion models in civil engineering.
keywords:
diffusion models, deep learning, computational design, building information modelling
@@ 1.1 Motivation
The new AI art generators show a remarkable capability in generating images of architectural design sketches and frontal or perspective views that display a higher degree of creativity, even if they are technically just recombining the training material. This opens the question if they can be used to more challenging tasks like generating floor plans or sectional drawings which are nowadays often manually drawn on paper or in CAD (Computer Aided Design) tools.
The process of planning constructions in civil engineering is still a very manual and time-consuming process. It possesses enormous potential for boosting productivity by 50 % to 60 %, which could result in an extra 1.6 trillion USD being added annually to the value of the construction industry barbosa2017reinventing . AI approaches are considered a core enable of this productivity boost PAN2021103517 ; pan2023integrating ; debrah2022artificial .
The recent breakthroughs of the AI art generation tools are based on the introduction of diffusion models in deep learning. They were first introduced by Sohl-Dickstein et al. in sohl2015deep and successively destroy information in an image during the training phase to learn how to recreate it. We will explain this process in more detail in Section 1.2.4.
The main topic of our paper is how these diffusion models can be used for computational design of architectural drawings like floors plans and how we can improve them to do this better. Within this paper we want to:
@@
  • Propose different semantic image-based diffusion model encodings
  • Compare the performance and workflows of those encodings
@@
  • Derive the requirements and architecture for a BIM-based diffusion model
@@
1.2 State of the Art
1.2.1 Parametric Design
@@
The idea of automating design steps with technology is an old one and various approaches developed over the time from parametric design to generative, and algorithmic design. Caetano, Santos, and Leitão combine them in their review under the term Computational Design caetano2020computational .
The idea of Parametric Design is that components are created and shaped according to parameterized functions in a computer, in contrast to being designed directly gursel2012creative ; jabi2013parametric ; schnabel2007parametric . It is founded in modelling shape attributes trough mathematical functions and was already used before computers were introduced e. g. by Antoni Gaudi in designing elements in the Sagrada Familia burry1993expiatory . Nowadays, the shape attributes are altered by computers which usually generates a large set of candidates that are then for example validated structurally monedero2000parametric ; rempling2019automatic or in their energy performancegranadeiro2013building to identify good ones to be presented to the user.
1.2.2 Generative Design
@@
Generative Design is using generative computer approaches like evolutionary or agent-based algorithms to generate and evaluate designs caldas2008generation ; krish2011practical . More recent approaches are starting to adopt Graph representations gan2022bim and Deep Neural Networks for generative tasks min2023floor ; rahbar2022architectural ; weber2022automated . The challenge in those approaches is that they are bound in their generation capability by the underlying generation rules and representation (e. g. genome encoding) and cannot in a creative way recombine them. Further they are prone to generate random results outside of the requested targets due to their probabilistic or non-deterministic nature of their search caetano2020computational ; chaszar2016generating .
1.2.3 Algorithmic Design
@@
Algorithmic design approaches usually describe computer programs that are used to generate using plans based on rules and logic implemented in code or code grammars humppi2016algorithm ; oxman2017thinking ; caetano2019integration . More recent approaches combine such a shape grammar with reinforcement learning to generate energy efficient designs mandow2020architectural . These approaches require understanding of the code to control the generation process, which makes them less applicable in practise.
1.2.4 Diffusion models
@@
Diffusion Models revolutionized the area of generating art with AI. They were first introduced in sohl2015deep and are inspired by non-equilibrium thermodynamics. Diffusion-models are based on a Deep Neural Network (DNN) architecture that does two things: First, it successively destroys the information in the image it is trained on by adding white noise in a so called forward diffusion step. In each step it trains a DNN to be able to recover the lost information in a reverse diffusion step. The resulting image now contains some of the original information like style and colour and some newly added information which is associated to tags describing the image. As the DNN is only trained to repair a small piece of lost information in the image, the generation approach becomes more stable and solve a main challenge with traditional approaches that try to generate the whole image from scratch and used to fail in the process.
This destroy and repair process is then repeated at different levels of resolution. This removes the limitation that the DNN can only repair lost information as we can now create new images by first creating them at a very low resolution and then successively repairing them until we have a high-resolution image. It also allows the DNN for the different resolutions to specialize on different visual aspects. For example, lower resolutions layers learn things like image composition, middle layers learn object shapes, and high resolutions learn object surface materials and colours.
@@
This basic idea of diffusion was then used to create neural network architectures that are capable of using reverse diffusion (denoising) to create high-quality image outputs from noisy inputs ho2020denoising . The concept is nowadays widely applied beyond image generation in natural language processing, video and 3D model generation, pose detection, timeseries modelling, etc. yang2023diffusion . An example of this step-by-step process is shown in Fig. 1. Configuring these models to create the desired outputs used to be a process that required expert knowledge. More recent architecture variants like OpenAI’s GLIDE model nichol2021glide contain an encoder, which can take an arbitrary text prompt by a user and create a valid text encoding that can be fed into the connected diffusion model. This architecture also includes a second model, which upsamples the result of the diffusion model.
Most current commercial models usually use the CLIP (Contrastive Language Image Pre-training) architecture presented in radford2021learning and ramesh2021zero . CLIP is responsible for training the encoder and determining how the text encodings are linked to image parts in the diffusion model. From this point, every platform contains slight differences in model and encoder architecture. The specific encoder used there is called unClip, which includes an image encoder and encodes both text and image inputs into a joint representation space from which the diffusion model can create an image Ramesh22 .
@@
1.2.5 Fine-tuning Diffusion Models
@@
Fortunately, the new generation of diffusion models allows to refine these federated models with a few examples to a new domain, which is called fine-tuning or few-shot-learning. There are different approaches for doing that. One approach is to retrain the whole diffusion-model on the new dataset and refine the weights in the deep neural network (DNN) Ruiz_2023_CVPR . This creates a variant of the initial federated diffusion-model that is of the same size as the original network and quite computational expensive to refine all the weights on the DNN.
To avoid this, other approaches introduce new layers into the federated diffusion model and modify only their weights. This results in smaller and faster to train models as the original federated model is not modified and fewer weights need to be fine-tuned hu2021lora ; roich2022pivotal . The additional layers also enable the models to also learn new capabilities, like transforming outline sketches into full images zhang2023adding .
The final approach to fine-tune is not to modify the DNN, but, the input vector from the prompt. The text prompt send to the diffusion model needs to be encoded as number vector to be used as input layer for the DNN. The finetuning changes the weights of this word embedding vector. A process called textual inversion gal2022textual . The benefit of this approach is that it is very fast to train and results in very small fine-tuned model that only contains the modified word embedding vector.
1.2.6 3D Diffusion models
[…]
2 Diffusion models for Floor Plans
2.1 Integrated workflow
One goal of the paper is to understand how generative AI tools can support practitioners in their typical design workflows. For this we want to compare the workflows for the different diffusion models. For this comparison we want first want to discuss a idealized workflows would like that fully integrates and unlocks the potential of generative AI tools in the design process in our opinion. This workflow is illustrated in Fig. 2.
Figure 2: Idealized integrated workflow for AI-based design
In this workflow, a planner uses a text interface or speech to text interface to specify the keywords describing his targeted design. As alternative input the user may be able to sketch out a rough plan to provide e. g. the general shape of a more complex planning tasks. The AI system would refine these initial inputs by generating questions to collect additional requirements and clarify uncertainties in a dialog with the user. These requirements may encompass functional (e. g. room types, size, etc.), non-functional (e. g. materials), stylistic (e. g. shape and colours), or computational constraints (e. g. structural dynamics, energy-efficiency). It would then generate multiple design options using a federated diffusion model, validate and rank them and allow the user to select the best ones and edit areas that require more work. The final design should be exported to a BIM model like IFC that the user can easily import in his tools for further use. This workflow is based on the common workflows for diffusion models ec3 . We consider it also good for civil engineering, because the user has multiple ways to control the design process from specify and refining requirements to editing model parts.
Within the next sections we will first discuss qualitatively, what elements of this workflow are already feasible with the existing AI tools and then evaluate their performance in different experiments in section 3. From this we derive further research topics needed to realize this idealized workflow in section 4.
2.2 Out-of-the-box Diffusion Model $\mathcal{B}$
Many users are adopting the new AI art tools to generate architectural drawings. The study in ec3 has shown that already 7 % of all prompts on the Midjourney platform are related to architecture with 14.741 prompts in mean per day. Of those prompts 43 % are related to floor plans.
Midjourney can be used with the Discord messaging app as it runs on the servers of the provider. This on the one hand makes it easy to use, as no local model needs to be installed and configured. On the other hand, all queries are run on the companies servers and by default are visible to other users. This poses a privacy and data security risk.
Stable Diffusion is a alternative generative AI solution that can be installed and operated locally. This overhead brings the benefit to the users to execute their prompts in private and to retrain the model for their specific applications.
@@ 2.2 Out-of-the-box Diffusion Model $\mathcal{B}$
The out-of-the-box approach for a user testing the technology is to use the out-of-the box model to generate his floor plans. We call this approach our baseline approach $\mathcal{B}$. Fig. 3 show the workflow for this. The user writes a text query describing the design they would like to have and sends it to the generative AI. Out-of-the-box, the generative AI uses a federated Diffusion model and generates one or multiple binary images as response to the query.
@@
The workflow of this out-of-the-box approach has multiple differences in comparison to Fig. 2. We do not have the refinement loops like that the AI model generates refining questions or the section editing parts of the generated design as they are not fully supported by the current technologies ploennigs2022ai . Further, the out-of-the-box approach creates only a binary image of a floor plan instead of directly generating a BIM model. In theory, this binary image could be used to automatically reconstruct a BIM model. However, this is a non-trivial problem still in research gimenez2016automatic ; zeng2019deep ; pizarro2022automatic . In practise, the user needs to manually create the BIM model from on the generated image.
@@
2.3 Fine-tuned Diffusion Model $\mathcal{R}$
The out-of-the-box approach $\mathcal{B}$ is prone to creating faulty results, as we will confirm in the experimental section 3. Fig. 4 shows an example floor plan generated by Stable Diffusion for the query “building floor plan of a one family house with a garden”. The image shows some resemblance to a floor plan, but also many incorrect elements. The reason for the low-quality result is that the out-of-the-box model is trained on a wide variety of images and will respond with images with a large variance and not the expected narrow focus on floor plans.
@@ 2.3 Fine-tuned Diffusion Model $\mathcal{R}$
Therefore, it is necessary to fine-tune the diffusion model to correctly understand what kind of floor plans we expect. This can be done by retraining the model with new sample images and queries that resemble the expected results as discussed in section 1.2.5. This modifies our workflow as illustrated in Fig. 5. We now have a training and a generation phase. In the training phase we fine-tune an existing federated diffusion model with a small set of sample images and queries. We use the tuned diffusion model in the generation phase to generate then designs that are closer to the training data.
@@
2.4 Fine-tuned Diffusion Models with Semantic Encoded rooms $\mathcal{SR}$ and elements $\mathcal{SE}$
@@
Figure 6: Example generation results of the four retrained styles.
In previous work ec3 it was shown that one of the main reasons why diffusion models struggle with floor plans is the missing semantic understanding of the elements in the image. While a diffusion model may learn that the term ”floor plan” is correlated to rectangular lines, it does not really understand that these lines are walls enclosing rooms. The moment we add more symbols to the image like doors, windows, furniture or assets, the model will struggle understanding the meaning of these symbols and mix them up as visible in Fig. 4. The question is: Can make it easier for the diffusion model to learn these semantics?
We propose a novel approach to encode these semantics in floor plans to improve the fine-tuning of diffusion models utilizing the following three strategies:
Simplify:
The more different semantics a floor plan contains the harder it is for the diffusion model to differentiate. Thus, it is important to remove unnecessary elements (furniture, assets, etc.) and focus on the relevant semantics for floor planning (walls, doors, and windows). The removed unnecessary elements may be added in a following in-painting step in the section editing step from Fig. 2. This will be our baseline for the retraining of model $\mathcal{B}$, which will be called $\mathcal{R}$ (Reduced).
@@
When we look at a floor plan, we know that thick lines are walls and the enclosed white space is a room. We also understand its semantic function (kitchen, bed room, living room) from the symbols of appliances that it contains (oven/fridge, bed, couch). When we remove those symbols—to simplify the model—the diffusion model (and we) cannot understand this semantic. So, we need a way to re-encode this information, for example by filling the rooms with different colours like blue for kitchen, orange for a bed room, or green for a living room. We call this semantic colouring of rooms $\mathcal{SR}$.
@@
Floor plans are often black and white (BW) and diffusion models are designed to encode colour images. Thus, we lose lots of information capacity by just training them on black and white images. We already used colours to encode room semantics in the last step. The question is if we can further simplify the floor plans by replacing the symbols for doors and windows with colours instead. We call this semantic colouring of elements $\mathcal{SE}$. The combined semantic encoding of room and elements is the $\mathcal{SRE}$ model.
2.5 Fine-tuning Methodology

### v2

# Automating Computational Design with Generative AI

@@ Abstract
AI image generators based on diffusion models have recently garnered attention for their capability to create images from simple text prompts. However, for practical use in civil engineering they need to be able to create specific construction plans for given constraints. This paper investigates the potential of current AI generators in addressing such challenges, specifically for the creation of simple floor plans. We explain how the underlying diffusion-models work and propose novel refinement approaches to improve semantic encoding and generation quality. In several experiments we show that we can improve validity of generated floor plans from 6 % to 90 %. Based on these results we derive future research challenges considering building information modelling. With this we provide: (i) evaluation of current generative AIs; (ii) propose improved refinement approaches; (iii) evaluate them on various examples; (iv) derive future directions for diffusion models in civil engineering.
Keywords generative AI  $\cdot$ diffusion models $\cdot$ deep learning $\cdot$ computational design $\cdot$ building information modelling
@@ 1 Introduction
The process of planning for construction projects in civil engineering is still very manual and time-consuming, with architects and engineers spending numerous hours creating designs and plans [3]. There is enormous potential for boosting productivity by 50 % to 60 %, which could result in an extra 1.6 trillion USD being added annually to the value of the construction industry [4]. AI approaches are considered a core enabler of this productivity boost [5, 6, 7].
The recent breakthroughs of generative AI tools for images are based on the introduction of diffusion models in deep learning. They were first introduced by Sohl-Dickstein et al. in [8] and are nowadays the main approach for image and video generation (see Section 3.1). In this paper, we want to investigate how these diffusion models can be used for computational design of architectural drawings like floors plans and how we can improve their performance for this task. To achieve this, we do the following:
@@
  • Propose the fine-tuning of diffusion models with domain examples
  • Propose a novel semantic encoding of floor plans for improved fine-tuning
  • Fine-tune a model for each proposed encoding
  • Compare the performance and workflows for those encodings
@@
  • Derive the requirements and sketch the architecture for a BIM-based diffusion model
@@
We use the openly available, very commonly used Stable Diffusion v2.1 model as our basis for fine-tuning, to be able to realistically evaluate the capabilities of current diffusion models. To improve its results, we then propose a novel approach to fine-tune the diffusion models with floor plans rendered in specific visual styles to enable semantic encodings. Based on our experiments, we also identify the limitations of current approaches and sketch out the option of a BIM-based diffusion model as a potential solution, without implementing it, as this is beyond the scope of the paper and thus future work.
2 State of the Art in Computational Design
@@
The idea of automating design steps with technology is an old one. Various approaches developed over time, from Parametric to Generative, over Algorithmic Design. Caetano et al. grouped them the under the term Computational Design [9].
The idea of Parametric Design is that the parameters defining the shape (e. g. length, height, …) of components are varied by functions in a computer, in contrast to being set directly [10, 11, 12]. It was already used before computers were introduced e. g. by Antoni Gaudi for the Sagrada Familia [13]. Parametric design is used to generate variants of structural elements that are then validated and ranked structurally [14, 15] or in their energy performance[16].
@@
Generative Design is using evolutionary or agent-based algorithms to generate whole designs [17, 18]. The challenge in those approaches is that they are bound in their generation capability by the underlying generation rules and representations (e. g. genome encoding) and cannot recombine them in creative ways. Further they are prone to generate random results outside of the requested targets due to their probabilistic or non-deterministic nature of their search [9, 19].
@@
Algorithmic design approaches usually use rules or code grammars to create designs [20, 21, 22]. Recent approaches combine such a shape grammar with reinforcement learning to generate energy efficient designs [23]. These approaches are deterministic, but, require understanding of the rules/code to control the generation process, which makes them less applicable in practise.
A common use case across all approaches is the generation of floor plans[24]. Recently, also Generative Adversarial Networks, a Deep Neural Network architecture, are used to generate floor plans [25, 26, 27]. They are similar to diffusion models, which we will introduce in the next section.
3 Introduction to Diffusion models
3.1 Model Structure
@@
Diffusion Models revolutionized the area of generating art with AI. They were first introduced in [8] and are inspired by nonequilibrium thermodynamics. Diffusion-models are based on a Deep Neural Network (DNN) architecture that does two things: First, it successively destroys the information in the image it is trained on by adding white noise in a so called forward diffusion step. In each step it trains a DNN to be able to recover the lost information in a reverse diffusion step. The resulting image now contains some of the original information, like style and colour, and some newly added information which is associated to tokens describing the image. As the DNN is only trained to repair a small piece of lost information in the image, the generation approach becomes more stable and thus solves a major challenge with traditional approaches, which try to generate the whole image from scratch and often fail in attaining decent results.
Formally, they are formulated using a Markov chain of $T$ steps, meaning that each step only depends on the previous one. Given a pixel $x_{0}$ of an original image with distribution $q(x_{0})$, we define a forward diffusion process by adding Gausian noise with variance $\beta_{t}$ to $x_{t-1}$ with mean $\mu_{t}$ derived from $q(x_{t-1})$. This produces a new latent variable $x_{t}$ with distribution
$q(x_{t}|x_{t-1})=N(x_{t};\mu_{t}=\sqrt{1-\beta_{t}}x_{t-1};\beta_{t}).$ (1)
The trick now is, that if the steps, i. e. the variance $\beta_{t}$, are small enough, then the reverse diffusion can be assumed to be a Gaussian process for which the variance $\hat{\beta}_{\theta}$ depends only on $\beta_{t}$ and on a mean function $\hat{\mu}$ as follows
$p(x_{t-1}|x_{t})=N(x_{t-1};\hat{\mu}(x_{t},x_{0}|P);\hat{\beta}_{\theta}).$ (2)
This mean function $\hat{\mu}(x_{t},x_{0}|P)$ is now learned by a Deep Neural Network (DNN), into which we also feed the vector encoding $P$ of the tokens in a prompt describing the image[28].
@@
The recent generation of Diffusion Models is using the CLIP (Contrastive Language Image Pre-training) architecture presented in [29] and [30]. It is not simply conditioning the mean function $\hat{\mu}(x_{t},x_{0}|P)$ on the tokens vector $P$ describing the image, but learns how tokens are linked to image parts in the diffusion model. This additional abstraction layer allows CLIP to learn object types and image composition on its higher layer, and details such as surface materials and colours on the deeper diffusion layer. An example of this step-by-step process is shown in Fig. 1.
3.2 Fine-tuning Diffusion Models
@@
4 Methodologies for Generative Design
4.1 Idealized Integrated Workflow
One goal of this paper is to investigate how generative AI tools can support practitioners in their typical design workflows. In this section, we want to discuss an idealized workflow that, in our opinion, fully exploits the potential of generative AI tools in the design process. We will then also discuss different approaches to making aspects of this workflow feasible with currently existing models. For each of these approaches we will then describe the training process for a simple floor plan generation model in section 5 and quantitatively evaluate these models in section 6.
Figure 2: Idealized integrated workflow for AI-based design
In this workflow shown in Figure 2, a planner uses a text interface or speech to text interface to specify the keywords describing their targeted design. As an alternative input, the planner can sketch out a rough plan to provide e. g. the general shape of a more complex planning tasks. The AI system would refine these initial inputs by generating questions to collect additional requirements and clarify uncertainties in a dialog with the user. These requirements may encompass functional (e. g. room types, size, etc.), non-functional (e. g. materials), stylistic (e. g. shape and colours), or computational constraints (e. g. structural dynamics, energy-efficiency). It would then generate multiple design options using a federated diffusion model, validate and rank them and allow the user to select the best variants and edit areas that require more work. The final design should be exported to a BIM model like IFC, which users can easily import into their tools for further use. This workflow is based on common workflows for image diffusion models in the AI art domain [1]. We consider it a good fit for civil engineering, as it allows a user to to control the design process in multiple ways, from specifying and refining requirements, to editing model parts.
4.2 Existing Out-of-the-Box Workflow (Model $\mathcal{B}$)
@@
The most immediate way a user testing the technology will engage with it, is to run one of the pre-trained, out-of-the box models to generate their floor plans. We call this approach our baseline approach $\mathcal{B}$, which is described in Fig. 3. The user writes a text query describing their design idea and sends it to the generative AI.The generative AI then uses a federated diffusion model and generates one or multiple binary images as a response to the query.
@@
This workflow is different to our idealized workflow in Fig. 2 in multiple ways. There is no refinement loop that generates clarifying questions or allows us to edit sections, as they are not currently supported by the default models [1]. Further, the out-of-the-box approach creates only a binary image of a floor plan instead of directly generating a BIM model. In theory, this binary image could be used to automatically reconstruct a BIM model. However, this is a non-trivial problem with ongoing research [35, 36, 37]. In practise, the user still needs to manually create BIM models from the generated images.
@@
4.3 Proposed Fine-tuning Approach (Model $\mathcal{R}$)
To improve the generated results, we propose to fine-tune the out-of-the-box diffusion model with training images that follow a unified visual style and training prompts that accurately represent semantic properties such as the shown number and type of rooms. The out-of-the-box approach $\mathcal{B}$ is prone to creating faulty results, as we will confirm in the experimental section 6. Fig. 4 shows an example floor plan generated by Stable Diffusion v2.1 for the query “building floor plan of a one family house with a garden”. The image shows some resemblance to a floor plan, but also many incorrect elements. The reason for the low-quality result is that the out-of-the-box model is trained on a wide variety of images and will thus contain a large amount of visual variance instead of the expected narrow focus on floor plans.
@@ 4.3 Proposed Fine-tuning Approach (Model $\mathcal{R}$)
Therefore, it is necessary to fine-tune the diffusion model to train it to understand what specific kinds of floor plans are expected. This can be done by retraining the model with new sample images and queries that resemble the expected results as discussed in section 3.2. For this step, we utilize the first of three strategies we will describe throughout this section:
Simplify:
The more different semantics a floor plan contains, the harder it is for the diffusion model to differentiate between them. Thus, it is important to remove unnecessary elements (furniture, assets, etc.) and focus on the semantics relevant for floor planning (walls, doors, and windows). The removed unnecessary elements may be added in a following in-painting step in the section editing step from Fig. 2. This strategy will be our baseline for the retraining of model $\mathcal{B}$, which will be called $\mathcal{R}$ (Reduced).
This modifies our workflow as illustrated in Fig. 5. We now have a training and a generation phase. In the training phase we fine-tune an existing federated diffusion model with a small set of sample images and queries. We then use the tuned diffusion model in the generation phase to generate designs that are closer to the training data.
@@
4.4 Proposed Improved Fine-Tuning with Color Encoding of Rooms (Model $\mathcal{SR}$) and Elements (Model $\mathcal{SE}$)
@@
Figure 6: An example floor plan created by the procedural generation algorithm, rendered in each of the four encoding styles. Images like these are combined with a description prompt to be used as the training data for the four fine-tuning approaches.
To further improve the fine-tuning of the diffusion models, we propose a new visual encoding approach of semantic information in floor plans. Our experiments will show that one of the main reasons why diffusion models struggle with floor plans is the missing semantic understanding of the elements in the image. While a diffusion model may learn that the term "floor plan" is correlated to rectangular lines, it does not really understand that these lines are walls enclosing rooms. The moment we add more symbols to the image like doors, windows, furniture or assets, the model will struggle understanding the meaning of these symbols and mix them up as visible in Fig. 4. The question is: Can we make it easier for the diffusion model to learn these semantics?
To achieve this, we propose a novel approach to encode these semantics in floor plans to improve the fine-tuning of diffusion models by adopting the following two strategy in addition to Simplify:
@@
When we look at a floor plan, we know that thick lines are walls and the enclosed white space is a room. We also understand its semantic function (kitchen, bedroom, living room) from the symbols of appliances that it contains (oven/fridge, bed, couch). When we remove those symbols—to simplify the model—the diffusion model (and we) cannot gain understanding about these semantics from the image. Thus, we need a way to re-encode this information, for example by filling the rooms with different colours, like cyan for bathrooms, yellow for bed/living rooms, or magenta for kitchens. We call this semantic colouring of rooms $\mathcal{SR}$.
@@
Floor plans are often black and white (BW) and diffusion models are designed to encode colour images. Thus, we lose lots of information capacity by just training them on black and white images. We already used colours to encode room semantics in the last step. The question then is, whether we can further simplify the floor plans by replacing the symbols for doors and windows with colours instead. We call this semantic colouring of elements $\mathcal{SE}$. The combined semantic encoding of room and elements is the $\mathcal{SRE}$ model.
5 Model Training
5.1 Generation of the Training Data